If Mary and Carter can paint a house in 5 hours together and Mary can do the same job in 8 hours alone, how long would it take for Carter working alone?
Answer by nerdybill(7384)

If Mary and Carter can paint a house in 5 hours together and Mary can do the same job in 8 hours alone, how long would it take for Carter working alone?
.
Let x = time it takes for Carter to work alone
then
5(1/x + 1/8) = 1
5/x + 5/8 = 1
Multiplying both sides by 8x:
40 + 5x = 8x
40 = 3x
40/3 = x
13 1/3 hours = x
or
13 hours 20 minutes